# Rescuing 331 bileptons from the Landau pole

### v1

@@ Abstract
Among the particles being searched for at the LHC beyond the Standard Model are bileptons, which are doubly charged gauge bosons. Bileptons are predicted by several Standard Model extensions, including the so-called 331 models with $\beta=\sqrt{3}$. The minimal formulation of these models is generally plagued by a "low energy" Landau pole, which can undermine current predictions for TeV-scale bileptons. We analyze this issue and investigate the possibility to shift the Landau pole at higher energy scales by extending the troublesome minimal 331 models.
1 Introduction
@@
From the phenomenological point of view, 331 models predict the existence of new fermions and gauge bosons beyond the Standard Model (BSM). The gauge properties of the specific 331 model depend on an arbitrary parameter, denoted as $\beta$. In the case $\beta=\sqrt{3}$, one extra gauge boson is neutral and two have charges $Q=\pm 1$; we call them $Z^{\prime},{W^{\prime}}^{\pm}$, respectively. The remaining two extra gauge bosons are doubly charged and are denoted as $Y^{\pm\pm}$. Since they couple with two standard leptons they are called bileptons. From the perspective of collider physics, the presence of such a bilepton is one of the most interesting features of these 331 models. The existence of bileptons has been searched at LHC and recent limits on their masses have been set [5, 6, 7, 8]. The downside of 331 models with $\beta=\sqrt{3}$ is the well-known presence of a Landau pole at low-energy scales. Already in the 1950s Landau realized that the QED running coupling diverges—developing a pole—at a finite momentum scale. The presence of such a divergence, known as the Landau pole, is not limited to QED but poses a potential problem for a generic quantum field theory in the perturbative regime. The modern perspective on this issue involves several arguments. One key point is that the Landau pole arises within perturbation theory, which cannot be expected to reliably represent behaviors related to large or even infinite couplings. Real insight can only be gained using non-perturbative approaches. It is not unreasonable to assume that physical observables of a quantum field theory are well-behaved even when the coupling diverges at the Landau pole and beyond. This possibility has been explored and several models have been identified where this behavior is realized (see e.g. [9] and references therein). Another perspective emerges when we examine the energy scales of the Landau pole. In the Standard Model the Landau pole is above the Planck scale, but still so high that it becomes of no practical significance to study the SM in that regime. This leads to view theories with a Landau pole as UV-incomplete or effective theories. The presence of a Landau pole is taken as an evidence of an energy scale at which the theory breaks down and new physics must intervene. This interpretation becomes obviously more concerning when the scale of the Landau pole approaches energies accessible to present or near future experimental efforts. Since the running of the couplings changes according to the matter content of the models, it is necessary to investigate SM extensions for the presence of a Landau pole.
In the case of 331 models, the scale of the Landau pole depends on the value of the parameter $\beta$. Minimal 331 extensions with $\beta=1/\sqrt{3}$ exhibit a Landau pole in the $U(1)$ coupling at very high energies (much like the Standard Model), while setting $\beta=\sqrt{3}$ lowers the Landau pole to about $4$ TeV. Such a low energy divergence can be troublesome for bileptons detection in a near future by LHC. It is interesting both from the phenomenological and theoretical point of view to investigate how to build extensions of the minimal 331 model with $\beta=\sqrt{3}$ where the Landau pole occurs at higher energy scales [10, 11, 12, 13].
In this paper we address the Landau pole problem considering next-to-minimal 331 extensions. We find that by extending the model with a fourth family of quarks and leptons or by enlarging the Higgs sector, it is possible to shift the Landau pole at higher scales.
2 General features of 331 models
@@
    The charges of the first two components of triplet (or antitriplet) must match the charges of the corresponding SM fields. This matching fixes uniquely the $X$-values for the SM fermions as functions of $\beta$. The non-standard fields have the same $X$-values of the other members of the respective (anti)triplet, because of gauge invariance. Hence each quark (or lepton) family in a $\boldsymbol{3}$ representation shares the same $X$-value and the same happens for each quark (or lepton) family in a $\boldsymbol{\bar{3}}$ representation. This constraint does not need to apply when we add triplets (or antitriplets) composed solely of left-handed BSM fermions to the matter content. However, by assumption, we choose to retain this constraint. Both standard and non-standard left-handed quarks need a right-handed components for the model to be $SU(3)_{C}$ vector-like. This does not apply to leptons. However, for the anomaly cancellation and in analogy with the SM, we assume that each left-handed non-standard charged lepton has a right-handed counterpart of the same electric charge. Neutral leptons do not participate to anomaly cancellation, therefore we do not need to assume a right-handed counterpart.
@@
In the case $|\beta|=1/\sqrt{3}$ the BSM quarks have SM charges $-1/3$ and $2/3$; there are also 3 BSM neutral bosons (including $Z^{\prime}$) and charged BSM boson $W^{{}^{\prime}\pm}$. In the case $|\beta|=\sqrt{3}$ the extra quarks have non SM charges (4/3 and 5/3). The charged BSM bosons $W^{{}^{\prime}\pm}$ are accompanied by double charged bosons $Y^{\pm\pm}$, that can decay into a pair of same charge SM leptons, giving an interesting signature for LHC [2, 3, 5, 8]. They are indicated as bileptons, because they do not couple at tree level to SM quark pairs. Therefore the case $|\beta|=\sqrt{3}$ presents more exotic features compared to the case $|\beta|=1/\sqrt{3}$ and seems more distinctive for LHC detection.
3 Minimal 331-model with $\beta=\sqrt{3}$
@@
The $SU_{L}(3)$ triplet (or anti-triplet) has SM fermions occupying the first two components and a BSM fermion the third one. Since according to (9) the lepton multiplets transform as $\bar{3}$ under $SU_{L}(3)$, they take the general form $(\ell_{L},-\nu_{L}^{\ell},\xi^{\ell}_{L})$, where the index $\ell$ runs on the 3 families and $\xi^{\ell}_{L}$ is a left-handed BSM lepton. We note that each $\xi^{\ell}_{L}$ has the same electric charge of $\ell_{R}^{c}$ ($c$ stands for conjugate), namely $Q(\xi^{\ell}_{L})=+1$, as well as the same $X$-value, and it belongs to the same 331 representation. The economical identification $\xi^{\ell}_{L}\equiv\ell_{R}^{c}$ is the assumption which characterizes the so-called minimal 331 model. Let us observe that we cannot repeat the assumption in the quark sector, since all the non-standard extra quark fields have unconventional electric charges ($-4/3$ or $5/3$). The matter content of the minimal 331 model is summarized in Table 2. The gauge symmetry of 331 models is spontaneously broken in two steps as
$SU_{L}(3)\times U_{X}(1)\xrightarrow[]{\mu_{331}}SU_{L}(2)\times U_{Y}(1)\xrightarrow[]{\mu_{\rm EW}}U_{\rm e.m.}(1)$ (10)
@@
giving masses to the five extra gauge bosons. Among them, the bileptons $Y^{\pm\pm}$, doubly charged gauge bosons which are a distinctive feature of this model. The mass of the bilepton $Y$ depends both on the coupling $u$ and $SU(2)_{L}$ coupling $g_{2L}$, since [4]
@@
A lower bound of $\sim 1.3$ TeV at $90\%$C.L. [8] can be inferred by experimental constraints (see e.g. [17]). By using the value [18] $g_{2L}(M_{Z})=\sqrt{4\pi\alpha_{2L}(M_{Z})}\simeq 0.65$ one finds
@@
where the VEVs are related with the SM one $v_{EW}$ as $v^{2}+v^{\prime 2}=v_{EW}^{2}=(246\,\mathrm{GeV})^{2}\equiv\mu_{EW}^{2}$. Since the EW symmetry breaking should occur at a lower scale, we expect the hierarchy $u\gg v,v^{\prime},v_{EW}$. At this passage, three other generators are broken, which leads us to the three massive SM gauge bosons and one massless gauge boson (the photon).
@@
3.1 The Landau Pole in the minimal 331 model
The gauge coupling $g_{i}$ associated to the gauge groups $U(1),\,SU(2),\,SU(3)$ in (10), or equivalentely the couplings $\alpha_{i}\equiv g_{i}^{2}/4\pi$, are not constant and their values depend on the energy scale $\mu$. In perturbation theory, the expression for the running couplings $\alpha_{i}(\mu)$ given by the renormalization group in the so-called leading logarithm approximation is
@@ 3.1 The Landau Pole in the minimal 331 model
$b_{i}=\frac{2}{3}\sum_{fermions}\text{Tr}(T_{a}T_{a})+\frac{1}{3}\sum_{scalars}\text{Tr}(T_{a}T_{a})-\frac{11}{3}C_{2},$ (17)
where $C_{2}$ and $T_{a}$ are respectively the Casimir and the generators of the gauge group. The value of $b_{i}$, and therefore the scale of the Landau pole, critically depend on the matter content of the model. More precisely, the dependence is on how many fermions and scalars are active at the energy under scrutiny. Since for the $U(1)$ coupling $C_{2}=0$, the $b$ coefficient is positive, hence it always exhibits a Landau pole. In the case of the SM gauge group $SU_{L}(2)\times U_{Y}(1)\times SU_{c}(3)$ we have:
@@
$b_{Y}=41/6\qquad b_{2L}=-19/6\qquad b_{c}=-7\quad$ ${\rm SM}$ (19)
$b_{Y}=7\qquad b_{2L}=-3\qquad\qquad b_{c}=-7\quad$ ${\rm 2HDM}$ (20)
@@
In all 331 models the two scales $\mu_{331}$ and $\mu_{\mathrm{E}W}$ delimitate energy regimes with different symmetries and matter content, both of which affect the running. The SM gauge group dictates the running behavior of the gauge couplings $g_{Y},\,g_{2L}$ from the electroweak scale $\mu_{\rm EW}\sim M_{Z}$ up to the $\mu_{331}$ scale.
Above $\mu_{311}$ the heavy degrees of freedom become on-shell. The different matter content and the unbroken 331 gauge symmetry lead the running of $g_{X},\,g_{3L}$ above the $\mu_{331}$ scale. The matching condition at $\mu_{331}$ is given by [4]
$g_{2L}(\mu_{331})=g_{3L}(\mu_{331})\,,\qquad\frac{1}{g_{X}^{2}(\mu_{331})}=\left.\frac{1}{6}\left(\frac{1}{g_{Y}^{2}}-\beta^{2}\frac{1}{g_{L}^{2}}\right)\right|_{\mu=\mu_{331}}.$ (21)
@@
$\alpha_{Y}^{-1}(M_{Z})\simeq 98.341\,,\qquad\alpha_{2L}^{-1}(M_{Z})\simeq 29.589\,,\qquad\alpha_{c}^{-1}(M_{Z})\simeq 8.42\,.$ (22)
@@
Figure 1: We report the dependence of $\alpha_{X}^{-1}$ on the energy scale $\mu_{331}$. The dot-dashed and continuous lines correspond respectively to the cases where one and two Higgs doublets are on-shell below $\mu_{331}$.The vertical dashed line represents the Landau pole at $3908$ GeV.
To determine the running behavior of $\alpha_{X}^{-1}$, and hence its dependency on the matching scale $\mu_{331}$, we need to clearly define the matter content both above and $\mu_{331}$. In the scalar sector we expect that at least one of the three Higgs doublets has mass $\simeq\mu_{331}$ and contributes to the running only above the $\mu_{331}$ scale. We assume that the heavy Higgs, the exotic gauge fields $W^{\prime},Z^{\prime},Y$ and the exotic fermions $D^{1},\,D^{2},\,T$ have masses $\simeq\mu_{331}$, hence they are not active below that scale. One of the other Higgs doublets can be identified with the SM Higgs, leaving two alternatives for the remaining Higgs doublet:
@@
In figure (1) we show $\alpha_{X}^{-1}(\mu_{331})$ in case i), dot-dashed line, and ii), continuous line. A Landau pole arises when $\alpha_{X}^{-1}(\mu_{331})$ becomes zero, which occurs at at $\mu_{331}=3620$ GeV (case i)) and at $\mu_{331}=3908$ GeV (case ii)). These value set the upper limit of the 331 breaking scale $\mu^{\rm max}_{331}$. We have:
@@
Given the lower bound (14), namely $\mu_{331}\gtrsim 3850$   GeV, case i) is excluded by experimental data. We are left with case ii), which assumes two Higgs of mass lower than $\mu_{331}$.
The scale is constrained in the narrow range
@@
According to eqs. (13) and (14), the bilepton mass cannot exceed $\mu^{\rm max}_{331}$. Let us observe that, by definition of the mixing angle $\theta_{W}$, we can write
@@
Figure 2: (Left) Running of $\alpha_{i}^{-1}$ with the energy scale $\mu$. The vertical dashed line is the arbitrary 331 breaking scale $\mu_{331}$. (Right) Zoom of $\alpha_{X}^{-1}$. We compare with the multi-Higgs case with $N_{\rho}=4$ (dashed line) considered in section 4.
@@
The mass values of bileptons, given by $m_{Y}\sim\mu_{331}$, is close to the Landau pole. There is a narrowing gap between the areas where current experiments can look for bileptons and where the perturbation theory ceases to be accurate. This renders the minimal 331 model with $\beta=\sqrt{3}$ unreliable.
@@
4 Extending the scalar sector

### v2

@@ Abstract
Among the particles being searched for at the LHC beyond the Standard Model are bileptons, which are doubly charged gauge bosons. Bileptons are predicted by several Standard Model extensions, including the so-called 331 models with $\beta=\sqrt{3}$. The minimal formulation of these models is generally plagued by a low-energy Landau pole, which can undermine current predictions for TeV-scale bileptons. We analyze this issue and investigate the possibility to shift the Landau pole at higher energy scales by extending the troublesome minimal 331 models.
I Introduction
@@
From the phenomenological point of view, 331 models predict the existence of new fermions and gauge bosons beyond the Standard Model (BSM). The gauge properties of the specific 331 model depend on an arbitrary parameter, denoted as $\beta$. In the case $\beta=\sqrt{3}$, one extra gauge boson is neutral and two have charges $Q=\pm 1$; we call them $Z^{\prime},{W^{\prime}}^{\pm}$, respectively. The remaining two extra gauge bosons are doubly charged and are denoted as $Y^{\pm\pm}$. Since they couple with two standard leptons they are called bileptons. From the perspective of collider physics, the presence of such a bilepton is one of the most interesting features of these 331 models. The existence of bileptons has been actively investigated at LHC, bringing about lower bounds on their masses [5, 6]. Several phenomenological studies have explored their possible signatures at LHC (e.g. [7, 8]). The downside of 331 models with $\beta=\sqrt{3}$ is the well-known presence of a Landau pole at low-energy scales. Already in the 1950s Landau realized that the QED running coupling diverges—developing a pole—at a finite momentum scale. The presence of such a divergence, known as the Landau pole, is not limited to QED but poses a potential problem for a generic quantum field theory in the perturbative regime. The modern perspective on this issue involves several arguments. One key point is that the Landau pole arises within perturbation theory, which cannot be expected to reliably represent behaviors related to large or even infinite couplings. Real insight can only be gained using non-perturbative approaches. It is not unreasonable to assume that physical observables of a quantum field theory are well-behaved even when the coupling diverges at the Landau pole and beyond. This possibility has been explored and several models have been identified where this behavior is realized (see e.g. [9] and references therein). Another perspective emerges when we examine the energy scales of the Landau pole. In the Standard Model the Landau pole is above the Planck scale, thus it becomes of no practical significance to study the SM in that regime. This leads to view theories with a Landau pole as UV-incomplete or effective theories. The presence of a Landau pole is taken as an evidence of an energy scale at which the theory breaks down and new physics must intervene. This interpretation becomes obviously more concerning when the scale of the Landau pole approaches energies accessible to present or near future experiments. Since the running of the couplings changes according to the matter content of the models, it is necessary to investigate SM extensions for the presence of a Landau pole. In the case of 331 models, the scale of the Landau pole depends on the value of the parameter $\beta$. Minimal 331 extensions with $\beta=1/\sqrt{3}$ exhibit a Landau pole in the $U(1)$ coupling at very high energies (much like the Standard Model), while setting $\beta=\sqrt{3}$ lowers the Landau pole to about 444 TeV. Such a low energy divergence can be troublesome for bileptons detection in a near future by LHC. It is interesting both from the phenomenological and theoretical point of view to investigate how to build extensions of the minimal 331 model with $\beta=\sqrt{3}$ where the Landau pole occurs at higher energy scales [10, 11, 12, 13].
In this paper we address the Landau pole problem considering next-to-minimal 331 extensions.
II General features of 331 models
@@
    The charges of the first two components of triplet (or antitriplet) must match the charges of the corresponding SM fields. This matching fixes uniquely the $X$-values for the SM fermions as functions of $\beta$. The non-standard fields have the same $X$-values of the other members of the respective (anti)triplet, because of gauge invariance. Hence each quark (or lepton) family in a $\boldsymbol{3}$ representation shares the same $X$-value and the same happens for each quark (or lepton) family in a $\boldsymbol{\bar{3}}$ representation. This constraint does not need to apply when we add triplets (or antitriplets) composed solely of left-handed BSM fermions to the matter content. However, by assumption, we choose to retain this constraint. In analogy to Standard Model, each non-standard left-handed quark needs a right-handed component to maintain $SU(3)_{C}$ vector-like couplings. Obviously, that does not apply to leptons. However, for the anomaly cancellation and in analogy with the SM, we assume that each left-handed non-standard charged lepton has a right-handed counterpart of the same electric charge. Neutral leptons do not participate to anomaly cancellation, therefore we do not need to assume a right-handed counterpart.
@@
In the case $|\beta|=1/\sqrt{3}$ the BSM quarks have SM charges 1/3-1/3- 1 / 3 and 2/32/32 / 3; there are also 3 BSM neutral bosons (including $Z^{\prime}$) and charged BSM boson $W^{{}^{\prime}\pm}$. In the case $|\beta|=\sqrt{3}$ the extra quarks have non SM charges (4/3 and 5/3). The charged BSM bosons $W^{{}^{\prime}\pm}$ are accompanied by double charged bosons $Y^{\pm\pm}$, that can decay into a pair of same charge SM leptons, giving an interesting signature for LHC [2, 3, 7, 6]. They are indicated as bileptons, because they do not couple at tree level to SM quark pairs. Therefore the case $|\beta|=\sqrt{3}$ presents more exotic features compared to the case $|\beta|=1/\sqrt{3}$ and seems more distinctive for LHC detection. In the following sections we will study extensions of the so-called minimal 331 model with $\beta=\sqrt{3}$
III Minimal 331-model
@@
The $SU_{L}(3)$ triplet (or anti-triplet) has SM fermions occupying the first two components and a BSM fermion the third one. Since according to (9) the lepton multiplets transform as $\bar{3}$ under $SU_{L}(3)$, they take the general form $(\ell_{L},-\nu_{L}^{\ell},\xi^{\ell}_{L})$, where the index $\ell$ runs on the 3 families and $\xi^{\ell}_{L}$ is a left-handed BSM lepton. We note that each $\xi^{\ell}_{L}$ has the same electric charge of $\ell_{R}^{c}$ ($c$ stands for conjugate), namely $Q(\xi^{\ell}_{L})=+1$, as well as the same $X$-value, and it belongs to the same 331 representation. The economical identification $\xi^{\ell}_{L}\equiv\ell_{R}^{c}$ is the assumption which characterizes the so-called minimal 331 model. Let us observe that we cannot repeat the assumption in the quark sector, since all BSM quark fields have unconventional electric charges (4/3-4/3- 4 / 3 or 5/35/35 / 3). The matter content of the minimal 331 model is summarized in Table 3. The gauge symmetry of 331 models is spontaneously broken in two steps, to the Standard Model and the electroweak symmetries, namely
$SU_{L}(3)\times U_{X}(1)\xrightarrow[]{\mu_{331}}SU_{L}(2)\times U_{Y}(1)$
$SU_{L}(2)\times U_{Y}(1)\xrightarrow[]{\mu_{\rm EW}}U_{\rm e.m.}$ (10)
@@
giving masses to the five BSM gauge bosons. Among them, the bileptons $Y^{\pm\pm}$, doubly charged gauge bosons, are a distinctive feature of this model. They are called bileptons because they carry lepton number $L=\pm 2$ and couple only to standard lepton pairs, not to standard quarks.
The mass of the bilepton $Y$ depends on both $u$ and the $SU(2)_{L}$ coupling $g_{2L}$, since [16]
@@
This relation is derived under the assumption that the VEV scale $u$ is much larger than the VEVs responsible of the SM symmetry breaking (see below).
[…]
By using the value [19] $g_{2L}(M_{Z})=\sqrt{4\pi\alpha_{2L}(M_{Z})}\simeq 0.65$ and assuming the lower bound $m_{Y}\gtrsim 1.3$ TeV one finds
@@
where the VEVs are related to the SM one $v_{EW}$ by $v^{2}+v^{\prime 2}=v_{EW}^{2}=(246\,\mathrm{GeV})^{2}\equiv\mu_{EW}^{2}$. Since the EW symmetry breaking should occur at a lower scale, we expect the hierarchy $u\gg v,v^{\prime},v_{EW}$. We summarize the scalar content of the minimal 331 model in table 1.
Scalar $SU_{L}(3)$ $U_{X}(1)$
$\chi$ 3 +1+1+ 1
$\rho$ 3 0
$\eta$ 3 1-1- 1
Table 1: Scalar content of minimal 331 model.
At this passage, three other generators are broken, which leads us to the three massive SM gauge bosons and one massless gauge boson (the photon).
@@
III.1 The Landau Pole in the minimal 331 model
The gauge coupling $g_{i}$ associated to the gauge groups $U(1),\,SU(2),\,SU(3)$, or equivalentely the couplings $\alpha_{i}\equiv g_{i}^{2}/4\pi$, are not constant and their values depend on the energy scale $\mu$. In perturbation theory, the expression for the running couplings $\alpha_{i}(\mu)$ given by the renormalization group in the so-called leading logarithm approximation is
@@ III.1 The Landau Pole in the minimal 331 model
$b_{i}$ $=\frac{2}{3}\sum_{fermions}\text{Tr}(T_{a}T_{a})+$
$+\frac{1}{3}\sum_{scalars}\text{Tr}(T_{a}T_{a})-\frac{11}{3}C_{2},$ (17)
where $C_{2}$ and $T_{a}$ are respectively the Casimir and the generators of the gauge group. The value of $b_{i}$, and therefore the scale of the Landau pole, critically depend on the matter content of the model. More precisely, the dependence is on how many fermions and scalars are active at the energy under scrutiny. Since $C_{2}=0$ for the $U(1)$ gauge group, the $b$ coefficient is positive, hence the $U(1)$ coupling always exhibits a Landau pole. In the case of the SM gauge group $SU_{L}(2)\times U_{Y}(1)\times SU_{c}(3)$ we have:
@@
$b_{Y}=\frac{41}{6}\quad b_{2L}=-\frac{19}{6}\quad b_{c}=-7\quad{\rm SM}$ (19)
$b_{Y}=7\quad b_{2L}=-3\quad b_{c}=-7\quad{\rm 2HDM}$ (20)
@@
In all 331 models the two scales $\mu_{331}$ and $\mu_{\mathrm{E}W}$ delimitate energy regimes with different symmetries and matter content, both of which affect the running.
The SM gauge group dictates the running behavior of the gauge couplings $g_{Y},\,g_{2L}$ from the electroweak scale $\mu_{\rm EW}\sim M_{Z}$ up to the $\mu_{331}$ scale. Above $\mu_{311}$ the heavy degrees of freedom become on-shell. The different matter content and the unbroken 331 gauge symmetry lead the running of $g_{X},\,g_{3L}$ above the $\mu_{331}$ scale. The matching condition at $\mu_{331}$ is given by [16]
$g_{2L}(\mu_{331})=g_{3L}(\mu_{331})\,,$
$\frac{1}{g_{X}^{2}(\mu_{331})}=\frac{1}{6}\left(\frac{1}{g_{Y}^{2}(\mu_{331})}-\beta^{2}\frac{1}{g_{L}^{2}(\mu_{331})}\right).$ (21)
@@
$\alpha_{Y}^{-1}(M_{Z})$ 98.341$\simeq$ 98.341\,,≃ 98.341 ,
$\alpha_{2L}^{-1}(M_{Z})$ 29.589$\simeq$ 29.589\,,≃ 29.589 , (22)
$\alpha_{c}^{-1}(M_{Z})$ 8.42$\simeq$ 8.42\,.≃ 8.42 .
@@
Figure 1: We report the dependence of $\alpha_{X}^{-1}$ on the energy scale $\mu_{331}$. The dot-dashed and continuous lines correspond respectively to the cases where one and two Higgs doublets are on-shell below $\mu_{331}$. The vertical dashed line represents the Landau pole at 390839083908 GeV.
To determine the running behavior of $\alpha_{X}^{-1}$, and hence its dependency on the matching scale $\mu_{331}$, we need to clearly define the matter content both above and below $\mu_{331}$. In the scalar sector we expect that at least one of the three Higgs doublets has mass $\simeq\mu_{331}$ and contributes to the running only above the $\mu_{331}$ scale. We assume that the heavy Higgs, the exotic gauge fields $W^{\prime},Z^{\prime},Y$ and the exotic fermions $D^{1},\,D^{2},\,T$ have masses $\simeq\mu_{331}$, hence they are not active below that scale. One of the other Higgs doublets can be identified with the SM Higgs, leaving two alternatives for the remaining Higgs doublet:
@@
In figure (1) we show $\alpha_{X}^{-1}(\mu_{331})$ in case i), dot-dashed line, and ii), continuous line. A Landau pole arises when $\alpha_{X}^{-1}(\mu_{331})$ becomes zero, which occurs at at $\mu_{331}=3620$ GeV (case i)) and at $\mu_{331}=3908$ GeV (case ii)). These value set the upper limit of the 331 breaking scale $\mu^{\rm max}_{331}$. Thus we have:
@@
Given the lower bound (14), namely $\mu_{331}\gtrsim 3850$   GeV, case i) is excluded by experimental data.
The only possibility remains case ii), which assumes two Higgs of mass lower than $\mu_{331}$. The scale $\mu_{331}$ is constrained in the narrow range
@@
According to eqs. (13) and (14), the bilepton mass cannot exceed $\mu^{\rm max}_{331}$.
At the upper limit $\mu^{\rm max}_{331}$, $\alpha_{X}^{-1}$ hits the Landau pole. Let us observe that, by definition of the mixing angle $\theta_{W}$, we can write
@@
Figure 2: (Above) Running of $\alpha_{i}^{-1}$ with the energy scale $\mu$. The vertical dashed line is the arbitrary 331 breaking scale $\mu_{331}$. (Below) Zoom of $\alpha_{X}^{-1}$. We compare with the multi-Higgs case with $N_{\rho}=4$ (dashed line) considered in section IV.
@@
The mass values of bileptons, given by $m_{Y}\sim\mu_{331}/3$, is close to the Landau pole. There is a narrowing gap between the areas where current experiments can look for bileptons and where the perturbation theory ceases to be accurate.
@@
IV Extending the scalar sector